OpenStudy (anonymous):

Find the slope of the line that contains the points named. A(3, 2), B(7, 8)

OpenStudy (zzr0ck3r):

the slope \(m\) of the line that contains the points named \((x_0,y_0),(x_1,y_1)\) is given by \(m=\frac{y_0-y_1}{x_0-x_1}\)

OpenStudy (zzr0ck3r):

or \(m = \frac{y_1-y_0}{x_1-x_0}\)....same thing

OpenStudy (anonymous):

not understanding this its frustrating

OpenStudy (zzr0ck3r):

your points are \((3,2)\) and \((7,8)\) I showed it for \((x_0,y_0)\) and \((x_1,y_1)\) can you see how \(x_0=3, y_0=2, x_1=7 \) and \(y_1=8\) ?

OpenStudy (anonymous):

ok got it thank you so much

OpenStudy (zzr0ck3r):

the reason I am using these weird variables, is so you know how to do it in general, for ANY two points

OpenStudy (zzr0ck3r):

so plugging those in the equation I gave you, what do you get?

OpenStudy (zzr0ck3r):

Slope = \(\frac{8-2}{7-3}=\frac{6}{4}=\frac{3}{2}\)

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (anonymous):

yes it does thank you so much

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

you just subtract one y value from the other y value and put that over the same thing with the corresponding x values. if you don't match up the corresponding x values the sign will be wrong

OpenStudy (anonymous):

so this one would be C(3, 8), D(-2, 5) -3/5

OpenStudy (zzr0ck3r):

\(\frac{5-8}{-2-3}=\frac{-3}{-5}=\frac{3}{5}=\frac{8-5}{3-(-2)}\)

OpenStudy (zzr0ck3r):

the y's must mach up (be right on top) of the corresponding x values

OpenStudy (anonymous):

ok3/5

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

ok I am going to try this know if I need help ill be back and thank you again

OpenStudy (zzr0ck3r):

np